# Normal forms for ordinary differential operators, I

### v3

@@ 2.4 Some necessary conditions on the Schur operator
The coefficient matrix is again in Vandermonde’s shape, so there are only 00 solutions, this means $\overrightarrow{\alpha}=0$. Hence $det(T)\neq 0$. ∎
@@
The proof is by induction on $t$. All claims are obvious for $S_{-1}$. Suppose they are true from $S_{-1}$ to $S_{-i+1}$. Consider the homogeneous decomposition of $Q$:
@@
Note that $Q_{j}$ is a HCP for any $j$, $Sdeg_{B}(Q_{j})=-\infty$ for any $j$ and $Sdeg_{A}(Q_{q-i})\leq i-2$ for any $i>1$.
Since $QS=S\partial^{q}$, we have the equality of their $(q-i)$ homogeneous components: $(QS)_{q-i}=(S\partial^{q})_{q-i}=S_{-i}\partial^{q}$, i.e.
$\partial^{q}\cdot S_{-i}+Q_{q-2}\cdot S_{-i+2}+\dots+Q_{q-i}-S_{-i}\partial^{q}=0.$ (2.14)
Put $M_{i}:=-(Q_{q-2}\cdot S_{-i+2}+\dots+Q_{q-i})$, so that this equation becomes
$[\partial^{q},S_{-i}]=M_{i}.$
Then by lemma 2.10 we get that $M_{i}$ is a HCP and $Sdeg_{A}(M_{i})\leq i-2$ (the only possible term of degree $i-2$ is $Q_{q-i}$:
$Q_{q-i}=(\eta_{0;q-i}+\eta_{1;q-i+1}\Gamma_{1}+\dots+\eta_{i-2;q-2}\Gamma_{i-2})\partial^{q-i},$
whose degree depends on vanishing the coefficient $\eta_{i-2;q-2}$). From formulae (2.9) we get also $Sdeg_{B}(M_{i})=-\infty$. Now applying Lemma 2.9, we get $Sdeg_{A}(S_{-i})\leq i-1$ and $Sdeg_{B}(S_{-i})=-\infty$ (note that since $S_{-i}$ is homogeneous, $S_{-i}$ is a HCP).
@@
$\partial^{pq}\cdot S_{-t}+(Q^{p})_{pq-2}\cdot S_{-t+2}+\dots+(Q^{p})_{pq-t}=S_{-t}\partial^{pq}.$
@@ 2.5 Some necessary conditions on normal forms
For a given pair of monic operators $Q,P\in D_{1}$ with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p\geq 0$ we define a normal form of $P$ with respect to $Q$ as the operator $P^{\prime}:=S^{-1}PS$, where $S$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$.
@@
The Schur operator is not uniquely defined, but up to multiplication by the elements of the centralizer $C(\partial^{q})\subset\hat{D}_{1}^{sym}$ from proposition 2.1. By this reason the normal form of the operator $P$ is not uniquely defined, but up to conjugation by elements from this centralizer.
@@
1$\Rightarrow$2: It is easy to see that any differential operator $P\in D_{1}$ with constant highest symbol, i.e. $P=a_{q}\partial^{q}+\sum_{i=1}^{q}a_{q-i}\partial^{q-i}$ with $a_{q}\in K$, satisfies condition $A_{p}(0)$ for any $p>0$ with an extra property: all homogeneous components $P_{i}$ don’t contain $A_{i}$, cf. lemma 2.6.
2$\Rightarrow$3 is obvious.
@@
we want to show $xH$ is totally free of $B_{j}$, i.e to show for any $k\in\mathbb{Z}$, $xHD^{k}$ doesn’t contain $B_{j}$. Since $H$ is totally free of $B_{j}$, we know $Sdeg_{B}(HD^{k})=-\infty$. According to Lemma 2.10 (notice $-\mathop{\rm\bf ord}(\int)=1$), we know
@@ 3 Normal forms for commuting operators
In this section we give a convenient description of the centraliser $C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ and of normal forms of the operator $P$ with respect to $Q$. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve (according to the well known classification theory of commuting ordinary differential operators such sheaves describe rank one subrings of commuting operators with a given spectral curve, cf. [47, Th. 10.26]).
Consider the subring $\tilde{K}[A_{1},\ldots,A_{k-1}]\subset\hat{D}_{1}^{sym}$. Clearly, $\tilde{K}[A_{1},\ldots,A_{k-1}]\cong\tilde{K}[x]/(x^{k}-1)$. Note that we have an isomorphism of $\tilde{K}$-algebras
$\Phi:\tilde{K}[A_{1},\ldots,A_{k-1}]\rightarrow\tilde{K}^{\oplus k},\quad P\mapsto(1\circ P,\ldots,\partial^{k-1}\circ P)$
(here $\partial^{l}\circ P$ is the notation from definition 2.2, i.e. if $P=\sum_{i=0}^{k-1}p_{i}A_{i}$, then $\partial^{l}\circ P=\sum p_{i}\xi^{il}$, and $\tilde{K}^{\oplus k}$ denotes the semisimple algebra - the direct sum of algebras $\tilde{K}$). Indeed, $\Phi$ is obviously linear, and, since
@@
The ring $E_{k}$ is endowed with a natural discrete pseudo-valuation, which we will denote as $-\mathop{\rm ord}\nolimits_{\tilde{D}}$ (i.e. $\mathop{\rm ord}\nolimits_{\tilde{D}}(\sum_{l=M}^{\infty}P_{l}\tilde{D}^{-l})=M$). We extend the usual terminology used in this paper also for operators from $E_{q}$ (such as the notion of the highest coefficient, monic operators, etc.)
@@
Next, note that any operator from the centralizer $C(\partial^{k})$ embedded into $\mathfrak{B}$ goes under the isomorphism $\psi:\mathfrak{B}\cong M_{k}(C(\mathfrak{B}))$ to a matrix with entries belonging to the subring $\tilde{K}[D^{k}]$. Indeed, we know from lemma 2.4 that all coefficients of homogeneous terms of negative order satisfy the following conditions: if $P\in\hat{\Phi}(C(\partial^{k}))\subset\mathfrak{B}$, and $p_{l}\in\tilde{K}^{\oplus k}$ denote its coefficients, then always $\tilde{p}_{l,j}=$ for $j=0,\ldots,-l-1$ if $l<0$. Therefore all homogeneous terms of negative order will go to matrices with constant coefficients (see the proof of lemma 3.1), and terms of non-negative order go to matrices with entries belonging to the subring $\tilde{K}[D^{k}]$. Finally, we can observe that in fact we get an isomorphism $C(\partial^{k})\simeq M_{k}(\tilde{K}[D^{k}])$.
@@
We will find the operator $S$ as the limit of a Cauchy sequence666with respect to the topology defined by the pseudo-valuation $-\mathop{\rm ord}\nolimits_{\tilde{D}}$. Assume
@@
This system is solvable and give explicit formulae for unknown variables $s_{l,j}$: for $r=1,\ldots,(m-1)$ from the first $(m-1)$ equations we get
@@
Normal forms of commuting differential operators can be normalised (s.t. the normalised normal form will be uniquely defined up to conjugation). If the orders of these operators are coprime, such a normalisation can be easily presented:
@@
where
$p_{l,j}=\begin{cases}\mbox{If $p\geq q$ then}\begin{cases}\mbox{$0$ for $j\in N_{p-l}$}\quad l>p-q\\ \mbox{$0$ for $j=0,\ldots,-l-1$}\quad l<0\end{cases}\\ \mbox{If $p<q$ then}\begin{cases}\mbox{$0$ for $j\in N_{p-l}$}\quad l\geq 0\\ \mbox{$0$ for $j\in N_{p-l}$ or $j=0,\ldots,-l-1$}\quad p-q<l<0\\ \mbox{$0$ for $j=0,\ldots,-l-1$}\quad l\leq p-q\end{cases}\end{cases}$
@@
$\begin{cases}p_{l,0}-\tilde{s}_{l,0}+\tilde{s}_{l,p}=\tilde{p}_{l,0}\\ p_{l,p}-\tilde{s}_{l,p}+\tilde{s}_{l,2p}=\tilde{p}_{l,1}\\ \ldots\\ p_{l,(q-1)p}-\tilde{s}_{l,(m-1)p}+\tilde{s}_{l,0}=\tilde{p}_{l,q-1}\end{cases}$
because $GCD(p,q)=1$ (here again all indices are taken modulo $q$). Now we have $\tilde{s}_{l,0}=0$ for all $l=1,\ldots,q-1$, and, starting with the first equation, we can see that for $j$-th equation, where $j\in N_{l}$, we can find $\tilde{s}_{l,jp}$ such that $\tilde{p}_{l,j-1}=0$. On the other hand, $\tilde{s}_{l,jp}=0$ for all $j\notin N_{l}$. Thus, we uniquely determine $\tilde{s}_{l}$ such that conditions of lemma satisfied for $l$-th homogeneous component of $\tilde{P}^{\prime\prime}$.
On the other hand, since $\tilde{P}^{\prime\prime}\in C(\partial^{q})$, we know from lemma 2.4 that always $\tilde{p}_{l,j}=0$ for $j=0,\ldots,-l-1$ if $l<0$ (in particular, the number of zero coefficients for $p-q<l<0$ in case $p<q$ is constant). Taking $S=\prod_{j=1}^{q-1}S_{j}$ we get the needed operator: the normal form $S^{-1}P^{\prime}S$ will satisfy all conditions as stated. This completes the proof of the first statement.
To prove the second statement (the uniqueness of such normalised normal form up to conjugation), first note that if $S\in C(\partial^{q})$ is invertible and $\mathop{\rm\bf ord}(S)=t>0$, then the coefficient of the HCP $S_{t}$ must be a zero divisor. Indeed, assume the converse. Suppose $\mathop{\rm\bf ord}(S^{-1})=r$. Then $(SS^{-1})_{t+r}\neq 0$ hence $r=-t<0$. From lemma 2.4 we know then the coefficient of the HCP $(S^{-1})_{r}$ is a zero divisor, hence the coefficient of the HCP $(S^{-1}S)_{t+r}=1$ must be a zero divisor, a contradiction. By the same reason there are no invertible $S$ with $\mathop{\rm\bf ord}(S)<0$ (because all its homogeneous components will be zero divisors by lemma 2.4).
@@
([47, Th. 10.26]) There is a one-to-one correspondence
$[B\subset{D_{1}}\mbox{\quad of rank $1$}]/\thicksim\longleftrightarrow[(C,p,{{\cal F}})\mbox{\quad of rank $1$}]/\simeq$
@@
  • $[B]$ means a class of equivalent commutative elliptic subrings (i.e. $B$ containing a monic differential operator), where $B\sim B^{\prime}$ iff $B=f^{-1}B^{\prime}f$, $f\in D_{1}^{*}=K[[x]]^{*}$.
@@
  • $C$ is a (irreducible and reduced) projective curve over $K$, $p$ is a regular $K$-point and ${{\cal F}}$ is a torsion free sheaf of rank one with $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$ (a spectral sheaf).
  • $\simeq$ means a natural isomorphism of triples.
@@
The self contained proof of this theorem in such a form is given in [47, Ch. 9, 10], and it uses two other correspondences between these data and equivalence classes of Schur pairs (we recall the definition below, as they will play a key role in our statements). The proof given in [47] is an elaborated version of Mulase’s proof from [30] in a spirit of works [33], [32] and their higher dimensional generalisations in [44], [21].
@@
Let $B\subset D_{1}$ be an elliptic subring. Then by Schur theory from [37], cf. [47, T. 4.6, C. 4.7], there exists an invertible operator $S=s_{0}+s_{1}\partial^{-1}+\ldots$ in the usual (Schur’s) ring of pseudo-differential operators $E=K[[x]]((\partial^{-1}))$ such that $A:=S^{-1}BS\subset K((\partial^{-1}))$. Consider the homomorphism (of vector spaces)
@@
For subrings in $K((z))$ we can introduce the same notion of rank as for subrings in $D_{1}$777here $z$ will play a role of $\partial^{-1}$ or $\tilde{D}^{-1}$:
@@
A self contained proof of this theorem see e.g. in [47, 10.3].
Let $C$ be a projective curve over $K$ and $p\in C$ be a regular $K$-point. Then $C_{0}:=C\backslash p$ is an affine curve. According to theorem 3.1 for any torsion free rank one sheaf ${{\cal F}}$ on $C$ there exists a normalised commutative elliptic ring of differential operators $B\subset D_{1}$ defined uniquely up to a scale transform, such that $B$ is isomorphic to the ring of regular functions on $C_{0}$, $B\simeq{{\cal O}}_{C_{0}}(C_{0})$. Vice versa, any such ring $B$ is isomorphic to the ring of regular functions on some affine curve $C_{0}$, which can be compactified with the help of one regular $K$-point.
@@
We’ll call an affine curve $C_{0}$ over $K$ as affine spectral curve if it can be compactified with the help of one regular $K$-point, i.e. if there exists a projective curve $C$ over $K$ and a regular $K$-point $p$ such that $C_{0}\simeq C\backslash p$ (note that such $C$ is uniquely defined up to an isomorphism, see e.g. [16, Ch1., §6]).
@@
defined for any coherent torsion free sheaf ${{\cal F}}$ and any trivialisations $\pi,\hat{\phi}$ (for details see [47, Ch.10]). If ${{\cal F}}={{\cal O}}_{C}$, then $W=A$ and the rank of $A$ is $1$. Moreover, elements of $A$ are Laurent series expansions of functions from ${{\cal O}}_{C}(C\backslash p)$, s.t. the orders of elements of $A$ are the pole orders of corresponding functions. Differential operators corresponding to elements of $A$ via the correspondence from theorems 3.1, 3.2 have the same order as the elements.
If $\mathop{\rm rk}({{\cal F}})=1$, then $\mathop{\rm rk}(A)=1$ and the data $(C,p,{{\cal F}},\pi)$ uniquely determines the class $[B]$ from theorem 3.1, see [47, Rem. 10.24].
@@
Let $B\subset D_{1}$ be an elliptic commutative subring of ODOs. Let $P_{1},\ldots,P_{m}$ be its monic generators over $K$ (any generators have constant highest coefficients, cf. [47, Ch.3]), such that $\mathop{\rm\bf ord}(P_{1})=q$ is coprime with the order of $P_{2}$.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal, $f_{i}\in K[T_{1},\ldots,T_{m}]$. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$. From lemma 3.4 we immediately get that the coefficients of all other normal forms $P_{i}^{\prime}$ are uniquely determined. Obviously, $f_{i}(P_{1}^{\prime},\ldots,P_{m}^{\prime})=0$ for all $i$, and therefore define a set of equations on the coefficients of $\hat{\Phi}(P_{i}^{\prime})\in\mathfrak{B}$ in an affine space. Note that any point of the affine algebraic set determined by these equations together with equations coming from commutation relations between $P_{2}^{\prime},\ldots,P_{m}^{\prime}$ defines a set of coefficients of commuting operators from $C(\partial^{q})$, and these operators generate a ring isomorphic to $B$888with the isomorphism sending them to $T_{1},\ldots,T_{m}$.
@@
We’ll call a commutative ring $B^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ generated over $\tilde{K}$ by monic operators $P_{1}^{\prime}=\partial^{q},P_{2}^{\prime},\ldots,P_{m}^{\prime}$, where $P_{2}^{\prime}$ is normalised and $q$ is coprime with the order of $P_{2}^{\prime}$, as a normalised normal form of rank one with respect to a (ordered) set of generators $P_{1}^{\prime},\ldots,P_{m}^{\prime}$, and we’ll call the coefficients of the operators $P_{i}^{\prime}$ as coordinates of $B^{\prime}$ (cf. lemma 3.4).

### v4

@@ 2.4 Some necessary conditions on the Schur operator
The coefficient matrix is again in Vandermonde’s shape, so there are only 0 solutions, this means $\overrightarrow{\alpha}=0$. Hence $det(T)\neq 0$. ∎
@@
The proof is by induction on $t$. All claims are obvious for $S_{-1}$. Suppose they are true from $S_{-1}$ to $S_{-t+1}$. Consider the homogeneous decomposition of $Q$:
@@
Note that $Q_{j}$ is a HCP for any $j$, $Sdeg_{B}(Q_{j})=$-$\infty$ for any $j$ and $Sdeg_{A}(Q_{q-t})\leq t-2$ for any $t>1$.
Since $QS=S\partial^{q}$, we have the equality of their (qt)(q-t) homogeneous components: $(QS)_{q-t}=(S\partial^{q})_{q-t}=S_{-t}\partial^{q}$, i.e.
$\partial^{q}S_{-t}+Q_{q-2}S_{-t+2}+\dots+Q_{q-t}-S_{-t}\partial^{q}=0.$ (2.14)
Put $M_{t}:=-(Q_{q-2}S_{-t+2}+\dots+Q_{q-t})$, so that this equation becomes
$\partial^{q},S_{-t}]=M_{t}.$
Then by lemma 2.10 we get that $M_{t}$ is a HCP and $Sdeg_{A}(M_{t})\leq t-2$ (the only possible term of degree t2t-2 is $Q_{q-t}$:
$Q_{q-t}=(\eta_{0;q-t}+\eta_{1;q-t+1}\Gamma_{1}+\dots+\eta_{t-2;q-2}\Gamma_{t-2})\partial^{q-t},$
whose degree depends on vanishing the coefficient $\eta_{t-2;q-2}$). From formulae (2.9) we get also $Sdeg_{B}(M_{t})=$-$\infty$. Now applying Lemma 2.9, we get $Sdeg_{A}(S_{-t})\leq t-1$ and $Sdeg_{B}(S_{-t})=$-$\infty$ (note that since $S_{-t}$ is homogeneous, $S_{-t}$ is a HCP).
@@
$\partial^{pq}S_{-t}+(Q^{p})_{pq-2}S_{-t+2}+\dots+(Q^{p})_{pq-t}=S_{-t}\partial^{pq}.$
@@ 2.5 Some necessary conditions on normal forms
For a given pair of monic operators $Q,P\in D_{1}$ with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p\geq 0$ we define a normal form of $P$ with respect to $Q$ as the operator $P^{\prime}:=S^{-1}PS$, where $S$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$, $\mathop{\rm\bf ord}(S)=0$, $S_{0}=1$, $S_{-1}=0$.
@@
The Schur operator is not uniquely defined, but up to multiplication by the invertible elements of order zero from the centralizer $C(\partial^{q})\subset\hat{D}_{1}^{sym}$ from proposition 2.1. By this reason the normal form of the operator $P$ is not uniquely defined, but up to conjugation by such elements from this centralizer.
@@
[…]
@@
we want to show $xH$ is totally free of $B_{j}$, i.e to show for any $k\in\mathbb{Z}$, $xHD^{k}$ doesn’t contain $B_{j}$. Since $H$ is totally free of $B_{j}$, we know $Sdeg_{B}(HD^{k})=$-$\infty$. According to Lemma 2.10 (notice $\mathop{\rm\bf ord}(\int)=1$), we know
@@ 3 Normal forms for commuting operators
The famous Burchnall-Chaundy lemma ([7]) says that any two commuting differential operators $P,Q\in D_{1}:=K[[x]][\partial]$ are algebraically dependent. More precisely, if the orders $p,q$ of operators $P,Q$ are coprime888i.e. the rank of the ring $K[P,Q]$ is 1, see e.g. [50] for relevant definitions, in particular [50, Lemma 5.23] for a proof of the Burchnall-Chaundy lemma in general form. The statement about the form of polynomial follows easily from the proof., then there exists an irreducible polynomial $f(X,Y)$ of weighted degree $v_{p,q}(f)=pq$ of special form (here the weighted degree is defined as in Dixmier’s paper [14], cf. item 4 in the List of Notations below): $f(X,Y)=\alpha X^{q}\pm Y^{p}+\ldots$ (here $\ldots$ mean terms of lower weighted degree, $0\neq\alpha\in K$; in particular, for coprime $p$ and $q$ the polynomial $f$ is automatically irreducible), such that $f(P,Q)=0$. A similar result for commuting operators of rank $r$ was established in [45] (cf. [35, Th. 2.11]), in this case $q=\mathop{\rm ord}\nolimits(Q)/r$, $p=\mathop{\rm ord}\nolimits(P)/r$, and again $GCD(p,q)=1$.
In this section we give a convenient description of the centraliser $C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ and of normal forms of the operator $P$ with respect to $Q$. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve (according to the well known classification theory of commuting ordinary differential operators such sheaves describe rank one subrings of commuting operators with a given spectral curve, cf. [50, Th. 10.26]).
Consider the subring $\tilde{K}[A_{1},\ldots,A_{k-1}]\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$. Clearly, $\tilde{K}[A_{1},\ldots,A_{k-1}]\cong\tilde{K}[x]/(x^{k}-1)$. Note that we have an isomorphism of $\tilde{K}-$algebras
$\Phi:\tilde{K}[A_{1},\ldots,A_{k-1}]\rightarrow\tilde{K}^{\oplus k},\quad P\mapsto(1\circ P,\ldots,(\partial^{k-1}\circ P){\partial^{-k+1}})$
(here $\partial^{l}\circ P$ is the notation from definition 2.2, i.e. if $P=\sum_{i=0}^{k-1}p_{i}A_{i}$, then $\partial^{l}\circ P=(\sum p_{i}\xi^{il}){\partial^{l}}$, and $\tilde{K}^{\oplus k}$ denotes the semisimple algebra - the direct sum of algebras $\tilde{K}$). Indeed, $\Phi$ is obviously linear, and, since
@@
The ring $E_{k}$ is endowed with a natural discrete pseudo-valuation, which we will denote as $\mathop{\rm ord}\nolimits_{\tilde{D}}$ (i.e. $\mathop{\rm ord}\nolimits_{\tilde{D}}(\sum_{l=M}^{\infty}P_{l}\tilde{D}^{-l})=M$). We extend the usual terminology used in this paper also for operators from $E_{q}$ (such as the notion of the highest coefficient, monic operators, etc.)
@@
Next, note that any operator from the centralizer $C(\partial^{k})$ embedded into $\mathfrak{B}$ goes under the isomorphism $\psi:\mathfrak{B}\cong M_{k}(C(\mathfrak{B}))$ to a matrix with entries belonging to the subring $\tilde{K}[D^{k}]$. Indeed, we know from lemma 2.5 that all coefficients of homogeneous terms of negative order satisfy the following conditions: if $P\in\hat{\Phi}(C(\partial^{k}))\subset\mathfrak{B}$, and $p_{l}\in\tilde{K}^{\oplus k}$ denote its coefficients, then always $\tilde{p}_{l,j}=0$ for $j=0,\ldots,-l-1$ if $l<0$. Therefore all homogeneous terms of negative order will go to matrices with constant coefficients (see the proof of lemma 3.1), and terms of non-negative order go to matrices with entries belonging to the subring $\tilde{K}[D^{k}]$. Finally, we can observe that in fact we get an isomorphism $C(\partial^{k})\simeq M_{k}(\tilde{K}[D^{k}])$.
@@
We will find the operator $S$ as the limit of a Cauchy sequence111111with respect to the topology defined by the pseudo-valuation $\mathop{\rm ord}\nolimits_{\tilde{D}}$. Assume
@@
This system is solvable and give explicit formulae for unknown variables $s_{l,j}$: for $r=1,\ldots,(m-1)$ from the first (m1)(m-1) equations we get
@@
Normal forms of commuting differential operators can be normalised (s.t. the normalised normal form will be uniquely defined in its conjugacy class). If the orders of these operators are coprime, such a normalisation can be easily presented:
@@
where $p_{l,j}=0$ for indices $j$ satisfying the following rules (all indices are taken mod $q$):
[…]
@@
$\begin{cases}p_{l,0}-\tilde{s}_{l,0}+\tilde{s}_{l,p}=\tilde{p}_{l,0}\\ p_{l,p}-\tilde{s}_{l,p}+\tilde{s}_{l,2p}=\tilde{p}_{l,p}\\ \ldots\\ p_{l,(q-1)p}-\tilde{s}_{l,(m-1)p}+\tilde{s}_{l,0}=\tilde{p}_{l,(q-1)p}\end{cases}$
because $GCD(p,q)=1$ (here again all indices are taken modulo $q$). Now we have $\tilde{s}_{l,0}=0$ for all $l=1,\ldots,q-1$, and, starting with the first equation, we can see that for $j$-th equation, where $j\in N_{l}$, we can find $\tilde{s}_{l,jp}$ such that $\tilde{p}_{l,(j-1)p}=0$. On the other hand, $\tilde{s}_{l,jp}=0$ for all $j\notin N_{l}$. Thus, we uniquely determine $\tilde{s}_{l}$ such that conditions of lemma satisfied for (pl)(p-l)-th homogeneous component of $\tilde{P}^{\prime\prime}$.
On the other hand, since $\tilde{P}^{\prime\prime}\in C(\partial^{q})$, we know from lemma 2.5 that always $\tilde{p}_{l,j}=0$ for $j=0,\ldots,-l-1$ if $l<0$ (in particular, the number of zero coefficients for pq<l<0p-q<l<0 in case $p<q$ is constant). Taking $S=\prod_{j=1}^{q-1}S_{j}$ we get the needed operator: the normal form $S^{-1}P^{\prime}S$ will satisfy all conditions as stated. This completes the proof of the first statement.
To prove the second statement (the uniqueness of such normalised normal form up to conjugation), first note that if $S\in C(\partial^{q})$ is invertible and $\mathop{\rm\bf ord}(S)=t>0$, then the coefficient of the HCP $S_{t}$ must be a zero divisor. Indeed, assume the converse. Suppose $\mathop{\rm\bf ord}(S^{-1})=r$. Then $(SS^{-1})_{t+r}\neq 0$ hence r=t<0r=-t<0. From lemma 2.5 we know then the coefficient of the HCP $(S^{-1})_{r}$ is a zero divisor, hence the coefficient of the HCP $(S^{-1}S)_{t+r}=1$ must be a zero divisor, a contradiction. By the same reason there are no invertible $S$ with $\mathop{\rm\bf ord}(S)<0$ (because all its homogeneous components will be zero divisors by lemma 2.5).
@@
([50, Th. 10.26]) There is a one-to-one correspondence
$B\subset{D_{1}}\mbox{\qquad of rank $1$}]/\thicksim\longleftrightarrow[(C,p,{{\cal F}})\mbox{\qquad of rank $1$}]/\simeq$
@@
  • $[B]$ means a class of equivalent commutative elliptic subrings (i.e. $B$ containing a monic differential operator), where $B\cong B^{\prime}$ iff $B=f^{-1}B^{\prime}f$, $f\in D_{1}^{*}=K[[x]]^{*}$.
@@
[…]
  • $\simeq$ on the right hand side means an equivalence class of triples up to a natural isomorphism.
@@
The self contained proof of this theorem in such a form is given in [50, Ch. 9, 10], and it uses two other correspondences between these data and equivalence classes of Schur pairs (we recall the definition below, as they will play a key role in our statements). The proof given in [50] is an elaborated version of Mulase’s proof from [30] in a spirit of works [33], [32] and their higher dimensional generalisations in [47], [22]. In one direction, from the subring $B$ to geometric spectral data, the construction of this correspondence is the following: consider the Rees ring $\tilde{B}=\oplus_{i\geq 0}B_{i}s^{i}\subset B[s]$, where $B_{i}:=\{P\in B|\quad\mathop{\rm ord}\nolimits(P)\leq ir\}$. The curve $C$ is defined as $C:=\mathop{\rm Proj}\tilde{B}$. The ideal $(s)$ defines a smooth point $p$ on $C$. Analogously, the spectral module $F$ is endowed with a natural filtration given by the order function: $F_{i}:=\{f\in F|\quad\mathop{\rm ord}\nolimits(f)\leq i\}$. With this filtration we can associate a graded module and corresponding associated spectral sheaf:
$^{(-1)}\tilde{F}:=\oplus_{i\geq 0}F_{ir-1}s^{i},\quad{{\cal F}}:=\mathop{\rm Proj}{}^{(-1)}\tilde{F}$.
In other direction, from the spectral data to the ring $B$, one can use the Krichever map from remark 3.7 below to construct a corresponding Schur pair (see definition 3.3 below), and then theorem 3.2 to construct the ring $B$.
@@
Let $B\subset D_{1}$ be an elliptic subring. Then by Schur theory from [37], cf. [50, T. 4.6, C. 4.7], there exists an invertible operator $S=s_{0}+s_{1}\partial^{-1}+\ldots$ in the usual (Schur’s) ring of pseudo-differential operators $E=K[[x]]((\partial^{-1}))$ such that $A:=S^{-1}BS\subset K((\partial^{-1}))$. Consider the homomorphism (of vector spaces)
@@
For subrings in $K((z))$ we can introduce the same notion of rank as for subrings in $D_{1}$121212here $z$ will play a role of $\partial^{-1}$ or $\tilde{D}^{-1}$:
@@
A self contained proof of this theorem see e.g. in [50, 10.3]. In one direction, from $B$ to the Schur pair, the construction of this correspondence is described above. In other direction the Sato theorem is used (see [38], cf. [30, Appendix]) to find the corresponding Sato operator $S$ such that $W=F\cdot S$ and then $B=SAS^{-1}$.
[…]
@@
[…]
@@
[…]
If $\mathop{\rm rk}({{\cal F}})=1$, then $\mathop{\rm rk}(A)=1$ and the data $(C,p,{{\cal F}},\pi)$ uniquely determines the class $[B]$ from theorem 3.1, see [50, Rem. 10.24].
@@
[…]
Let $B\subset D_{1}$ be an elliptic commutative subring of ODOs. Let $P_{1},\ldots,P_{m}$ be its monic generators over $K$ (any generators have constant highest coefficients, cf. [50, Ch.3]), such that $\mathop{\rm\bf ord}(P_{1})=q$ is coprime with the order of $P_{2}$.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal, $f_{i}\in K[T_{1},\ldots,T_{m}]$. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$. From lemma 3.4 we immediately get that the coefficients of all other normal forms $P_{i}^{\prime}$ are uniquely determined. Obviously, $f_{i}(P_{1}^{\prime},\ldots,P_{m}^{\prime})=0$ for all $i$, and therefore define a set of equations on the coefficients of $\hat{\Phi}(P_{i}^{\prime})\in\mathfrak{B}$ in an affine space. Note that any point of the affine algebraic set determined by these equations together with equations coming from commutation relations between $P_{2}^{\prime},\ldots,P_{m}^{\prime}$ defines a set of coefficients of commuting operators from $C(\partial^{q})$, and these operators generate a ring isomorphic to BB131313with the isomorphism sending them to $T_{1},\ldots,T_{m}$.
@@
We’ll call a commutative ring $B^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}$ generated over $K$ by monic operators $P_{1}^{\prime}=\partial^{q},P_{2}^{\prime},\ldots,P_{m}^{\prime}$, where $P_{2}^{\prime}$ is normalised and $q$ is coprime with the order of $P_{2}^{\prime}$, as a normalised normal form of rank one with respect to a (ordered) set of generators $P_{1}^{\prime},\ldots,P_{m}^{\prime}$, and we’ll call the coefficients of the operators $P_{i}^{\prime}$ as coordinates of $B^{\prime}$ (cf. lemma 3.4).